# What is the nature of late infall streamers?

### v1

@@ Appendix B Low accretion rate
The fall-back streamers discussed in the main text emerge in a dense environment, that is, where $\dot{M}_{\mathrm{sys}}=${10}^{-8}\text{\,}\mathrm{M_{\Sun}}\text{\,}{\mathrm{yr}}^{-1}$\$. However, some stars may undergo considerably lower accretion rates, down to ${\sim}${10}^{-9}\text{\,}\mathrm{M_{\Sun}}\text{\,}{\mathrm{yr}}^{-1}$\$ (e.g., Herczeg et al. 2023). If less material is accreted onto the disk, related accretion signatures are fainter, and streamers may become undetectable. Therefore, we investigated a low accretion rate of ${10}^{-9}\text{\,}\mathrm{M_{\Sun}}\text{\,}{\mathrm{yr}}^{-1}$ in simulation number 6, with a setup that is otherwise identical to simulation number 2, where multiple fall-back streamers are visible.
We find that, while the emerging structures are qualitatively similar dynamically, the streamers can indeed no longer be detected, as shown in Fig. 13. In the peak brightness temperature map, only a single component of the streamer originating from the top right can be seen; all other streamers and their subcomponents can no longer be detected. Its emission is insignificant ($T_{\mathrm{b,peak}}<$5\text{\,}\mathrm{K}$\$) compared to the emission in all previous cases and compared to the disk emission, so that this system would not be detected to be moving through an environment with strong turbulence. Almost no scattered light emission can be seen at flux levels ${10}^{-3}\text{\,}\mathrm{mJy}\text{\,}{\mathrm{arcsec}}^{-2}$\$, and only a small region of the top right streamer can be seen in the moment 0 map with a faint signal of $60\text{\,}\mathrm{mJy}\text{\,}{\mathrm{arcsec}}^{-2}\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$. We therefore conclude that, in order for fall-back streamers to be a plausible formation scenario for observed streamers, the infall rate needs to be of the order of ${\sim}${10}^{-8}\text{\,}\mathrm{M_{\Sun}}\text{\,}{\mathrm{yr}}^{-1}$\$.

### v2

@@ Appendix B Low accretion rate
The fallback streamers discussed in the main text emerge in a dense environment, that is, where $\dot{M}_{\mathrm{sys}}=${10}^{-8}\text{\,}\mathrm{M_{\Sun}}\text{\,}{\mathrm{yr}}^{-1}$\$. However, some stars may undergo considerably lower accretion rates, down to ${\sim}${10}^{-9}\text{\,}\mathrm{M_{\Sun}}\text{\,}{\mathrm{yr}}^{-1}$\$ (e.g., Herczeg et al. 2023). If less material is accreted onto the disk, related accretion signatures are fainter, and streamers may become undetectable. Therefore, we investigated a low accretion rate of ${10}^{-9}\text{\,}\mathrm{M_{\Sun}}\text{\,}{\mathrm{yr}}^{-1}$ in simulation number 6, with a setup that is otherwise identical to simulation number 2, where multiple fallback streamers are visible.
We find that, while the emerging structures are qualitatively similar dynamically, the streamers can indeed no longer be detected, as shown in Fig. 13. In the peak brightness temperature map, only a single component of the streamer originating from the top right can be seen; all other streamers and their subcomponents can no longer be detected. Its emission is insignificant ($T_{\mathrm{b,peak}}<$5\text{\,}\mathrm{K}$\$) compared to the emission in all previous cases and compared to the disk emission, so that this system would not be detected to be moving through an environment with strong turbulence. Almost no scattered light emission can be seen at flux levels ${10}^{-3}\text{\,}\mathrm{mJy}\text{\,}{\mathrm{arcsec}}^{-2}$\$, and only a small region of the top right streamer can be seen in the moment 0 map with a faint signal of $60\text{\,}\mathrm{mJy}\text{\,}{\mathrm{arcsec}}^{-2}\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$. We therefore conclude that, in order for fallback streamers to be a plausible formation scenario for observed streamers, the infall rate needs to be of the order of ${\sim}${10}^{-8}\text{\,}\mathrm{M_{\Sun}}\text{\,}{\mathrm{yr}}^{-1}$\$.